OpenStudy (aravindg):

how to factorize 2sin^2x-3sin x-2

OpenStudy (anonymous):

assume sin x =t then it becomes 2t^2 - 3t-2=0 now you are familiar???

OpenStudy (aravindg):

hmm ... wel u see i hav a problem in factorizing polynomials .can u hlp?

OpenStudy (aravindg):

hw do u actually factorize polynomials like these??i mean whats the method or steps to be followed?

OpenStudy (anonymous):

2t^2 - 4t +t-2=0 2t(t-2)+1(t-2)=0 (2t+1)(t-2)=0

OpenStudy (aravindg):

i feel some polynomials are non-intutive to be factorized by looking

OpenStudy (turingtest):

it's called factoring by grouping

OpenStudy (aravindg):

means wel how do u get intution that which term is to be grouped?

OpenStudy (anonymous):

try to break the middle term in to two numbers such that sum is equal to the value of b and product is c in the quadratic eq. ax^2 + bx + c=0

OpenStudy (turingtest):

...of course, there is no substitute for practice when it comes to developing an intuition on such things

OpenStudy (aravindg):

wel what to do if it involves cubic expressions?

OpenStudy (anonymous):

@aravindG do you get it??? yes @turingtest these can be developed effectively through practice.

OpenStudy (turingtest):

that depends

OpenStudy (aravindg):

like x^3-4x^2+4x

OpenStudy (anonymous):

take x common

OpenStudy (turingtest):

you can factor an x out, then factor the remaining quadratic

OpenStudy (anonymous):

then the same way!!!

OpenStudy (aravindg):

oh wow can u xplain this method of factorizing a quadratic

OpenStudy (aravindg):

i mean sum in middle product in end

OpenStudy (anonymous):

if you add -4 and +1 you get -3 (only such combination possible) while their product is also -4 matching the question..so you can proceed..

OpenStudy (aravindg):

oh thx i had a prob in that part i hav one more qn pls help i will post separately